# CoMAL: Collaborative Multi-Agent Large Language Models for Mixed-Autonomy Traffic

### v1

@@ Abstract
The integration of autonomous vehicles into urban traffic has great potential to improve efficiency by reducing congestion and optimizing traffic flow systematically. In this paper, we introduce CoMAL (Collaborative Multi-Agent LLMs), a framework designed to address the mixed-autonomy traffic problem by collaboration among autonomous vehicles to optimize traffic flow. CoMAL is built upon large language models, operating in an interactive traffic simulation environment. It utilizes a Perception Module to observe surrounding agents and a Memory Module to store strategies for each agent. The overall workflow includes a Collaboration Module that encourages autonomous vehicles to discuss the effective strategy and allocate roles, a reasoning engine to determine optimal behaviors based on assigned roles, and an Execution Module that controls vehicle actions using a hybrid approach combining rule-based models. Experimental results demonstrate that CoMAL achieves superior performance on the Flow benchmark. Additionally, we evaluate the impact of different language models and compare our framework with reinforcement learning approaches. It highlights the strong cooperative capability of LLM agents and presents a promising solution to the mixed-autonomy traffic challenge. The code is available at https://github.com/Hyan-Yao/CoMAL
@@ 1 Introduction
Recently, there has been significant growth in end-to-end autonomous driving systems [1]. The integration of large language models (LLMs) [2, 3] enhances the ability to generalize to unseen traffic scenarios with embedded common-sense knowledge. These models [4, 5, 6] primarily focus on optimizing the performance of individual ego vehicles. However, it also matters to study the problem of mixed-autonomy traffic to better deploy autonomous vehicles in society [7]. In this article, we enable connected autonomous vehicles (CAVs) to collaborate with human-driven vehicles across the traffic network, aiming to optimize overall traffic flow and system-wide efficiency.
@@
Traffic dynamics are extremely complex and chaotic dynamical systems [8]. Pioneering researchers [9, 10] typically trade away the complexity of the model for desirable provable properties. To this end, they propose a series of hand-designed control rules [11, 12] to direct a fleet of vehicles to form a desired stable motion pattern. The experiments [13, 14] suggest that autonomous vehicles can enhance traffic throughput, which highlights the potential of mixed-autonomy systems. To model the complex interactions between autonomous and human-driven vehicles in mixed-autonomy traffic systems, simulation-based numerical methods like the Flow benchmark [7, 15] have proven to be effective tools. Flow, a deep reinforcement learning (RL) framework, enables the systematic design of RL tasks to improve overall traffic flow, including the control of autonomous vehicles and traffic signals. With its ability to simulate dynamic multi-agent environments, Flow has facilitated the rise of multi-agent RL, which has gained popularity for modeling emergent behaviors in mixed-autonomy traffic systems [16]. By allowing autonomous vehicles to learn cooperative strategies, multi-agent RL enables interaction with both human-driven and other autonomous vehicles, optimizing traffic flow. These RL-based models demonstrate strong performance in specific traffic scenarios, effectively learning control policies for coordinating mixed-autonomy traffic [17, 18]. However, human behaviors in real-world settings are highly varied and unpredictable, while road networks are more complex and diverse. Deep RL performs well in specific and controlled environments but struggles to generalize across different scenarios [19] due to its reliance on large amounts of specific training data. Meanwhile, it is difficult to understand the decision-making process of RL, which limits its interpretability.
@@
Rethinking human behavioral patterns, adolescents can learn to drive in just 20 hours and handle unfamiliar situations [20], while young children can spontaneously cooperate to enhance work efficiency [21]. Human decision-making and cooperation are inherently knowledge-driven, relying on common sense, verbal communication, and reasoning [22]. This is in contrast to data-driven methods that require vast amounts of training data and often struggle with generalization. Recent advancements in LLMs [23] offer promising knowledge-driven solutions for addressing the mixed-autonomy traffic problem. LLMs can emulate human-like knowledge, including the ability to make decisions, form agreements, analyze road situations, and collaborate in real-time. By leveraging this embodied human knowledge, multiple intelligent LLM agents with distinct roles and attributes can work together to handle complex tasks more efficiently [24], offering new avenues for improving traffic systems through cooperation and common-sense reasoning.
@@
Building upon these insights and recognizing the limitations of RL in generalizing to new traffic situations, we propose leveraging LLMs for their common-sense reasoning and adaptability. LLMs offer better generalization by drawing on embedded knowledge, allowing them to manage complex and unpredictable traffic dynamics more effectively. We developed an innovative framework named CoMAL specifically designed to address mixed-autonomy traffic challenges, as depicted in Figure 1. CoMAL comprises a simulation environment that enables the interaction of individual LLM agents and allows agents to collect perceptual data from the traffic system. Leveraging the stored experiences within the Memory Module, the agents participate in a brainstorming session in the Collaboration Module, where they allocate tasks and establish their specific roles. Subsequently, each agent develops driving plans according to its designated role through the Reason Engine. Then Reason Engine generates a rule-based driving planner grounded in the Intelligent Driver Model (IDM), which is then implemented in the Execution Module to compensate for the inherent limitations of LLMs in control performance.
Our primary contributions are as follows:
@@
    To the best of our knowledge, we are the first to integrate the collaborative capability of multi-agent language models in autonomous driving. The Collaboration Module effectively establishes a session for multiple agents to engage in brainstorming and task allocation.
@@
    We evaluate CoMAL on Flow benchmark in three classical traffic scenarios (Figure 1 (c)) and compare its performance against RL methods. Extensive experimental results demonstrate significant performance improvements in terms of average velocity and driving smoothness across various LLMs.
@@
    We conduct experiments using the latest LLM models, including GPT-4o-mini and Qwen-72B/32B/7B, showcasing the adaptability of CoMAL across a diverse range of LLMs and demonstrating its superior performance over state-of-the-art RL approaches in mixed-autonomy traffic scenarios.
@@ 2.1 Mixed-Autonomy Traffic
Reinforcement learning (RL) offers a more dynamic and adaptable solution. Benchmarks in RL [25] like Mujoco and the Arcade Learning Environment [26] provide systematic evaluation and comparison of algorithms. And especially for Mixed-Autonomy, benchmark Flow [7] proposes four traffic scenarios to illustrate distinct RL problems including shockwave minimization, inflow management, efficient merging, and intersection control. It evaluates and compares RL algorithms like Trust Region Policy Optimization (TRPO) [15, 27], Proximal Policy Optimization (PPO) [28], Evolutionary Strategies (ES) [29], and Augmented Random Search (ARS) [30] in traffic scenarios. [31] also explores the mixed-autonomy scenario in a multi-agent traffic signal control system [32].
@@ 2.2 Large Language Model-based Multi-Agents
The communication structure of LLM-based multi-agent systems varies across different studies to address specific challenges  [23, 34, 24]. For example, research [24] has explored both centralized and decentralized communication structures for LLM-based multi-agent systems. In traffic control, LLMs facilitate human-machine interaction and improve decision-making processes. The survey [23] discusses how LLMs can be integrated into multi-agent systems to enhance communication and coordination among agents in traffic environments. And the study DiLu [22] explores a knowledge-driven approach using LLMs, while LLMLight [35] utilizes LLMs as decision-making agents for traffic signal control. A pioneering model Open-TI [36] that integrates LLMs with external traffic analysis tools to perform comprehensive traffic simulations and task-specific operations. Their findings suggest that incorporating domain-specific knowledge into LLMs can significantly enhance the performance and reliability of multi-agent systems in complex scenarios. This involves not only improving the performance of individual agents but also fostering effective communication and coordination among the agents.
@@ 3.2.1 Collaboration Module
Communication Structure Here, we introduce a shared message pool to boost communication efficiency, as shown in Figure 3. This communication structure maintains a shared message pool where agents can publish messages and subscribe to the latest messages from one another. Agents will take turns to speak one at a time until the strategy is fully developed and each agent’s role is clearly defined.
@@ 3.2.2 Reason Engine
The hierarchical chain-of-thought process involves four critical components: role clarification, scene understanding, motion instruction, and planner generation. Initially, it is crucial to clarify the role and task of the ego vehicle within a collaborative context. Then the LLM is directed to focus on key information in the scenario, such as headway distance and lead vehicles. Based on the scenario analysis, the LLM is then prompted to provide motion instructions for the ego vehicle. Finally, each agent utilizes scenario analysis and motion instructions to generate a driving planner, parameterized by IDM model.
@@ 4.3 Metrics
Figure 4: Visualization of vehicle trajectories in Ring 0 setting. The ring road has a total length of 230 meters and contains 22 vehicles. Each line in the space-time diagrams shows the position of a specific vehicle over time. When a vehicle completes a full lap of the ring, its position resets to zero. Left: In the absence of automated vehicles, human-driven vehicles exhibit stop-and-go shockwaves due to inherent instability. Right: With three connected autonomous vehicles using the CoMAL framework, the unstable vehicles are stabilized.
4.4 Specification on Communication
@@ 4.5 Quantitative Results
4.6 Ablation Studies
We conducted a detailed analysis of the effectiveness of each component of CoMAL in the FE 1 and Merge 1 scenarios, as presented in Table 3.
Ablation on Perception
Ablation on Memory
The comparisons in the third and fifth rows of Table 3 illustrate the impact of the Memory Module, in which specific experiences are allocated for each scenario. In the absence of high-quality experiences, agents are more susceptible to errors in both discussion and reasoning.
@@ 4.6 Ablation Studies
The comparison presented in the fourth and fifth rows of Table 3, as well as in the first row, highlights the effectiveness of the Collaboration Module. The absence of a collaboration module causes all agents to adopt nearly identical strategies, which in turn leads to conflicts and duplicated efforts. The absence of collaboration results in performance that can be even worse than that of the simple rule-based model shown in the first row.
@@ 4.7 Discussion
Compared to RL method
We conducted experiments on FE and the Merge scenarios by comparing them to RL methods developed in [7]. The results are shown in Table 5. The Ring is not included in this experiment because the adopted RL benchmark doesn’t have the Ring scenario embedded. In the FE scenario, CoMAL demonstrates robust global collaboration, whereas multi-agent RL models struggle to differentiate roles, hindering effective cooperation. Consequently, CoMAL outperforms RL-based approaches. However, in the Merge scenario, CoMAL performs less effectively than RL, indicating that the collaboration is not global. This finding highlights the critical importance of cooperation in enhancing performance.
@@
Comparison of various LLM models
@@ 5 Conclusion
In this paper, we present CoMAL, an effective LLM-based multi-agent framework to address mixed-autonomy traffic challenges. By prompt-tuning LLMs with a hierarchical LLM-based planner, CoMAL is able to handle complex vehicle driving tasks towards a collaborative goal under mixed-autonomy traffic. The LLM agent does so by serving primarily as a high-level commander, coordinating with lower-level controllers to execute detailed operations. Extensive experiments demonstrate the significant potential of multi-agent systems driven by LLMs to make informed decisions and collaborate effectively in driving scenarios.
We also acknowledge the limitations of our current work and would like to point out several important future directions. First, we can extend the experimental current settings to more agents and improve their collaboration to see if there would be emerging behaviors formed like in RL methods. Second, our paper addresses a simplified use of LLMs alone, whereas the combination of RL with LLM might be helpful in improving the performance of LLMs. In the future, more sophisticated scenarios will be explored to test the possibility of LLM’s behavior in complex tasks.
@@ Acknowledgements
We thank OpenAI for providing us with API credits under the Researcher Access program.

### v2

@@ Abstract
The integration of autonomous vehicles into urban traffic has great potential to improve efficiency by reducing congestion and optimizing traffic flow systematically. In this paper, we introduce CoMAL (Collaborative Multi-Agent LLMs), a framework designed to address the mixed-autonomy traffic problem by collaboration among autonomous vehicles to optimize traffic flow. CoMAL is built upon large language models and operates in an interactive traffic simulation environment. Specifically, It utilizes a Perception Module to observe surrounding agents and a Memory Module to store strategies for each agent. The overall workflow includes a Collaboration Module that encourages autonomous vehicles to discuss the effective strategy and allocate roles, a reasoning engine to determine optimal behaviors based on assigned roles, and an Execution Module that controls vehicle actions using a hybrid approach combining rule-based models. Experimental results demonstrate that CoMAL achieves superior performance on the Flow benchmark. Additionally, we evaluate the impact of different language models and compare our framework with reinforcement learning approaches. It highlights the strong cooperative capability of LLM agents and presents a promising solution to the mixed-autonomy traffic challenge. The code is available at https://github.com/Hyan-Yao/CoMAL
@@ 1 Introduction
Recently, there has been significant growth in end-to-end autonomous driving systems [1]. The integration of large language models (LLMs) [2, 3] enhances the ability to generalize to unseen traffic scenarios with embedded common-sense knowledge. These models [4, 5, 6] primarily focus on optimizing the performance of individual ego vehicles. However, it also matters to study the problem of mixed-autonomy traffic to better deploy autonomous vehicles in society [7]. In mixed-autonomy traffic, connected autonomous vehicles (CAVs) are enabled to collaborate with human-driven vehicles across the traffic network, aiming to optimize overall traffic flow and system-wide efficiency.
@@
To optimize the overall dynamics in mixed-autonomy traffic, existing researchers propose hand-designed control rules [11, 12, 9, 10] to direct a fleet of vehicles to form a desired stable motion pattern. The experiments [13, 14] suggest that autonomous vehicles can enhance traffic throughput, which highlights the potential of mixed-autonomy systems. To model the complex interactions between autonomous and human-driven vehicles in mixed-autonomy traffic systems, multi-agent reinforcement learning (MARL) methods have proven to be effective in learning cooperative strategies for autonomous vehicles in benchmarking environments like the Flow benchmark [7, 15]. While RL-based models demonstrate strong performance in specific controlled mixed-autonomy scenarios [17, 18], they showed several weaknesses. Firstly, RL methods struggle to generalize across different scenarios [19]. This is because in mixed-autonomy traffic, the varied and unpredictable behaviors of human drivers, combining with diverse road networks, makes the training RL rely on large amounts of data. Additionally, it is difficult to understand the decision-making process of RL, which limits its interpretability.
@@
Rethinking human behavioral patterns, adolescents can learn to drive in just 20 hours and handle unfamiliar situations [20], while youngest can spontaneously cooperate to enhance work efficiency [21]. Human decision-making and cooperation are inherently knowledge-driven, relying on common sense, verbal communication, and reasoning procedures [22]. This is in contrast to data-driven methods that require vast amounts of training data and often struggle with generalization. Recent advancements in LLMs [23] offer promising knowledge-driven solutions for addressing the mixed-autonomy traffic problem. LLMs can emulate human-like knowledge, including the ability to make decisions, form agreements, analyze road situations, and collaborate in real-time. By leveraging this embodied human knowledge, multiple intelligent LLM agents with distinct roles and attributes can work together to handle complex tasks more efficiently [24], offering new avenues for improving traffic systems through cooperation and common-sense reasoning.
@@
Building upon these insights and recognizing the limitations of RL in generalizing to new traffic situations, we propose leveraging LLMs for their common-sense reasoning and adaptability. LLMs offer better generalization by drawing on embedded knowledge, allowing them to manage complex and unpredictable traffic dynamics more effectively. We developed an innovative framework named CoMAL specifically designed to address mixed-autonomy traffic challenges, as depicted in Figure 1. CoMAL comprises a simulation environment that enables the interaction of individual LLM agents and allows agents to collect perceptual data from the traffic system. Leveraging the stored experiences within the Memory Module, the agents participate in a brainstorming session in the Collaboration Module, where they allocate tasks and establish their specific roles. Subsequently, each agent develops driving plans according to its designated role through the Reason Engine. Then the Engine generates a rule-based driving planner grounded in the Intelligent Driver Model (IDM), which is then implemented in the Execution Module to compensate for the inherent limitations of LLMs in control performance. Our primary contributions are as follows:
@@
    To the best of our knowledge, this paper is the first to integrate the collaborative capability of multi-agent language models in autonomous driving. The Collaboration Module effectively establishes a session for multiple agents to engage in brainstorming and task allocation.
@@
    We evaluate CoMAL on Flow benchmark in three classical traffic scenarios as shown in Figure 1 (c) and compare its performance against RL methods, with experimental results demonstrating significant performance improvements on average velocity and driving smoothness across various LLMs.
@@
    We have conducted experiments using multiple LLM models, including GPT-4o-mini and Qwen-72B/32B/7B, showcasing the adaptability of CoMAL across a diverse range of LLMs.
@@ 2.1 Mixed-Autonomy Traffic
Reinforcement learning (RL) offers a more dynamic and adaptable solution. Benchmarks in RL [25] like Mujoco and the Arcade Learning Environment [26] provide systematic evaluation and comparison of algorithms. Especially for Mixed-Autonomy, benchmark Flow [7] proposes four traffic scenarios to illustrate distinct RL problems including shockwave minimization, inflow management, efficient merging, and intersection control. It evaluates and compares RL algorithms like Trust Region Policy Optimization (TRPO) [15, 27], Proximal Policy Optimization (PPO) [28], Evolutionary Strategies (ES) [29], and Augmented Random Search (ARS) [30] in traffic scenarios. [31] also explores the mixed-autonomy scenario in a multi-agent traffic signal control system [32]. However, none of the existing work explores the LLMs’ applicability in mixed-autonomy traffic problems.
@@ 2.2 Large Language Model-based Multi-Agents
The communication structure of LLM-based multi-agent systems varies across different studies to address specific challenges  [23, 34, 24]. For example, research [24] has explored both centralized and decentralized communication structures for LLM-based multi-agent systems. In traffic control, LLMs facilitate human-machine interaction and improve decision-making processes. The survey [23] discusses how LLMs can be integrated into multi-agent systems to enhance communication and coordination among agents in traffic environments. And the study DiLu [22] explores a knowledge-driven approach using LLMs, while LLMLight [35] utilizes LLMs as decision-making agents for traffic signal control. A pioneering model Open-TI [36] that integrates LLMs with external traffic analysis tools to perform comprehensive traffic simulations and task-specific operations. Their findings suggest that incorporating domain-specific knowledge into LLMs can significantly enhance the performance and reliability of multi-agent systems in complex scenarios. Even though there exists work about LLM agents for traffic problems, forming better cooperation among agents remains challenging, this paper proposes a novel way to solve the mixed-autonomy task by efficient multi-agent coordinations.
@@ 3.2.1 Collaboration Module
Communication Structure In this section, we introduce a shared message pool to boost communication efficiency, as shown in Figure 3. This communication structure maintains a shared message pool where agents can publish messages and subscribe to the latest messages from one another. Agents will take turns to speak one at a time until the strategy is fully developed and each agent’s role is clearly defined.
@@ 3.2.2 Reason Engine
The hierarchical chain-of-thought process involves four critical components: role clarification, scene understanding, motion instruction, and planner generation. Initially, it is crucial to clarify the role and task of the ego vehicle within a collaborative context. Then the LLM is directed to focus on key information in the scenario, such as headway distance and lead vehicles. Based on the scenario analysis, the LLM is then prompted to provide motion instructions for the ego vehicle. Finally, each agent utilizes scenario analysis and motion instructions to generate a driving planner, parameterized by IDM model [38].
@@ 4.4 Specification on Communication
Figure 4: Visualization of vehicle trajectories in Ring 0 setting. The ring road has a total length of 230 meters and contains 22 vehicles. Each line in the space-time diagrams shows the position of a specific vehicle over time, whose speed is indicated with different colors. When a vehicle completes a full lap of the ring, its position resets to zero. Left: In the absence of automated vehicles, human-driven vehicles exhibit stop-and-go shockwaves due to inherent instability. Right: With three connected autonomous vehicles using the CoMAL framework, the unstable vehicles are stabilized.
@@ 4.5 Quantitative Results
4.6 Ablation Studies
We further conducted the detailed ablation analysis of the effectiveness of each component of CoMAL in the Figure Eight (FE) 1 and Merge 1 scenarios, the results are presented in Table 3.
Ablation on Perception
Ablation on Memory
The comparisons in the third and fifth rows of Table 3 illustrate the impact of the Memory Module, in which specific experiences are allocated for each scenario. In the absence of high-quality experiences, agents are more susceptible to errors in both discussion and reasoning thus causing a performance decrease in the overall optimization.
@@ 4.6 Ablation Studies
The comparison presented in the fourth and fifth rows of Table 3, as well as in the first row, highlights the effectiveness of the Collaboration Module. The absence of this module causes all agents to adopt nearly identical strategies in our experiment, which in turn leads to conflicts and duplicated efforts in execution. The absence of such collaboration results in performance that can be even worse than that of the simple rule-based model shown in the first row.
@@ 4.7 Discussion
Comparison with RL methods:
We conducted experiments on FE and the Merge scenarios and compared with RL methods developed in [7]. The results are shown in Table 5. The Ring is not included in this experiment because the adopted RL benchmark doesn’t have the Ring scenario embedded. In the FE scenario, CoMAL demonstrates robust global collaboration, whereas multi-agent RL models struggle to differentiate roles, hindering effective cooperation. Consequently, CoMAL outperforms RL-based approaches. However, in the Merge scenario, CoMAL performs less effectively than RL, indicating that the collaboration is not global. This finding highlights the critical importance of cooperation in enhancing performance.
@@
Comparison of various LLM models:
@@ 5 Conclusion
In this paper, we present the CoMAL, an effective LLM-based multi-agent framework to address mixed-autonomy traffic challenges. By prompt-tuning LLMs with a hierarchical LLM-based planner, CoMAL is able to handle complex vehicle driving tasks towards a collaborative goal under mixed-autonomy traffic. The LLM agent does so by serving primarily as a high-level commander, coordinating with lower-level controllers to execute detailed operations. We hope this work can demonstrate the significant potential of multi-agent systems driven by LLMs to make informed decisions and collaborate effectively in mixed-autonomy scenarios.
Besides the great potential, we also acknowledge the limitations of our current work and would like to point out several important future directions. First, a possible extension would be increasing the experimental scale to more agents, and improving their collaboration to observe if emerging behaviors would be formed, just like in RL methods. Second, our paper addresses such mix-autonomy problem by the simplified use of LLMs alone, whereas the combination of RL with LLM might be helpful to take advantage of both: exploration and contextual information, and thus improve the performance on this task. In the future, more sophisticated scenarios will be explored to test the possibility of LLM’s behavior in complex tasks.
@@ Acknowledgements
The work was partially supported by NSF awards #2421839, NAIRR #240120. This work used AWS through the CloudBank project, which is supported by National Science Foundation grant #1925001. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies. We thank OpenAI for providing us with API credits under the Researcher Access program and Amazon Research Awards.